OpenStudy (anonymous):

obtain the parametric equations of y^2 = 4x. the answers are x= 4/t^2 and y=4/t. why is this??

OpenStudy (anonymous):

This is because t is the parameter since we write parametric form in terms of r he took in terms of t

OpenStudy (anonymous):

so..

OpenStudy (anonymous):

okay..

OpenStudy (anonymous):

aah this is

OpenStudy (ikram002p):

well do this find y(t) it well abear that \[\left( \frac{ 4 }{ t } \right)^2=4\left( \frac{ 4 }{ t^2}\right)\] \[\frac{ 16 }{ t^2 }=\frac{ 16 }{ t^2 }\rightarrow L.H.S=R.H.S\]

OpenStudy (anonymous):

what is LHS??

OpenStudy (ikram002p):

left hand side

OpenStudy (anonymous):

so is it always that we have to assume that y is t?

OpenStudy (ikram002p):

no , assume f(x) = y and find f(t).

OpenStudy (anonymous):

im so sorry i still find it so difficult. my prof wasn't able to explained this topic to us.. so for example x^2 + 2x - y = 0?

OpenStudy (ikram002p):

is x & y are still defind as you mintioned in the Q ?

OpenStudy (anonymous):

it wasn't mentioned.

OpenStudy (ikram002p):

\[x^2 +2x-y=0\rightarrow y=x^2+2x\] assume f(x)=y \[f(x)=x^2+2x\] to check find f(t) if \[x=\frac{ 4 }{ t^2 } and \rightarrow y=\frac{ 4 }{ t } \] \[\frac{ 4 }{ t }=\left( \frac{ 4 }{ t^2 } \right)^2 + 2\left( \frac{ 4 }{ t} \right)\]

OpenStudy (anonymous):

x^2 + 2x - y = 0, the answers are x= t-2 and y = t(t-2) the answers are mention in the book

OpenStudy (ikram002p):

but if you want to find the root of x^2 + 2x - y = 0? y=x^2+2x you must have Points satisfy the equation so if you have point use Use the derivation

OpenStudy (ikram002p):

so you wana have to check the answer or prove it !

OpenStudy (anonymous):

derivation. so its 2x + 2

OpenStudy (ikram002p):

yes .

OpenStudy (anonymous):

aah i cant get x= t-2. why is this happeniing to me